Boxes of maraconi are filled by machine and the net weight of a box has a normal distribution with mean 10oz and standard deviation 0.7oz
If 25 boxes are selected at random, what is the probability that the average of their weights is greater tna 10.35oz?

Answer by stanbon(48568)

Boxes of maraconi are filled by machine and the net weight of a box has a normal distribution with mean 10oz and standard deviation 0.7oz
If 25 boxes are selected at random, what is the probability that the average of their weights is greater than 10.35oz?
-------------------------
The distribution of means of samples of size 25 has a mean of 10 oz
and a standard deviation of 0.7/sqrt(25) = 0.7/5 = 0.14
--------------
z(10.35) = (10.35-10)/0.14 = 2.5
P(z>2.5) = 0.006
So, Prob(a sample of size 25 has a mean > 10.35) = 0.006

============
Cheers,
Stan H.